# Neutral phylogenetic models and their role in tree-based biodiversity measures

### v2

@@ 4 Feature diversity (FD) - a generalisation of PD
It turns out that FD can be represented by PD on some tree when the collection $\mathcal{C}_{\mathbb{F}}$ of subsets $X_{f}=\{x\in X:f\in F_{x}\}$ is nested (i.e., it satisfies the condition that $A,B\in\mathcal{C}_{\mathbb{F}}\Rightarrow A\cap B\in\{A,B,\emptyset\}$), by results in [42] (Proposition 1 and Theorem 1). Moreover, since a nested collection of subsets of $X$ corresponds to a subset of the clusters of a phylogenetic tree, it is possible to describe the evolution of each character in $\mathbb{F}_{X}$ on that tree by a single-gain (of the feature on one edge $e$) and no subsequent loss (on any descendant edge of $e$). A slightly stronger statement is the following:
@@ 4.1 A stochastic model of FD on fixed trees
Figure 4: A simple example to illustrate the gain and loss of features on a tree and the FOB process (from [43]). Left: New features arise (indicated by $+$) and disappear (indicated by --) along edges of the tree. For convenience, no features are present at time 00. Five features are present among the leaves of the tree at time $t$: $\{\alpha,\beta,\gamma,\delta,\epsilon\}$. Right: Extinction at the present (denoted by $\dagger$) results in the loss of three extant species, leaving just three features ($\alpha,\beta,\gamma$) remaining among the leaves of the resulting tree, Thus, the ratio of surviving features to total features is $3/5=0.6$.
@@
  • new features arise (on any edge) at rate $r>0$ per unit time, and
@@
This is a classical (constant-birth, linear-death) stochastic process for which the following results hold (see e.g. [44], p. 461). Consider a directed path in the tree $T$ of length $L$ (the sum of the $\ell_{e}$ values on the path). Conditional on there being $k$ features present at the start of this path, let $N_{L}^{(k)}$ denote the number of features present at the end of this path. For $k=0$, $N_{L}^{(0)}$ has a Poisson distribution with mean $\frac{r}{\nu}(1-e^{-\nu L})$, and this mean is asymptotic to $rL$ as $L\rightarrow 0$ and to $r/\nu$ as $L\rightarrow\infty.$ More generally, if there are $k$ features present at the start of the path, then $N_{L}^{(k)}$ is the sum of two independent random variables: one distributed as $N_{L}^{(0)}$, the other has a Gamma distribution with parameters $(k,\nu L)$. In particular, as $L\rightarrow\infty$, $N_{L}^{(k)}$ converges to a Poisson distribution with mean $r/\nu$.
@@ 4.2 FD in birth–death trees with extinction at the present
In this section, we are interested in determining the proportion of features at the tips of the tree that survive the FOB model. Let $\mathcal{T}_{t}^{s}$ denote the tree obtained from $\mathcal{T}_{t}$ by deleting all leaves at time $t$ that do not survive the FOB model.
@@
Now, the number of species at time $t$ in $\mathcal{T}_{t}^{s}$ that have a copy of a particular feature that arose at some fixed time $t_{0}\in(0,t)$ in $\mathcal{T}_{t}$ is described exactly by the birth–death process $X_{t-t_{0}}$ with birth rate $\lambda$, death rate $\theta=\mu+\nu$ and extinction at the present (with survival probability $s$ for each individual). It follows that as $t$ becomes large, it becomes increasingly certain that none of the species at time $t$ in $\mathcal{T}_{t}$ will contain a given feature $f$ present at $t_{0}$ if $\lambda\leq\mu+\nu$, whereas if $\lambda>\mu+\nu$, there is a positive limiting probability that $f$ will be present in the extant leaves of the $\mathcal{T}_{t}$ (even after pruning under the FOB model).
@@
It is easily verified that $\varphi_{FD}(s)$ is an increasing concave function that satisfies $1\geq\varphi_{FD}(s)\geq s$ for all $s$. This is similar to the earlier story for PD. However, in contrast to $\varphi_{PD}(s)$, there does not seem to be a simple explicit expression for $\varphi_{FD}(s)$ when $\nu>0$ (when $\nu=0$, $\varphi_{FD}=\varphi_{PD}$ and so is given by eqns. (6) and (7)). Nevertheless, $\varphi_{FD}(s)$ can be expressed as a ratio of two integrals as follows. We will assume that $\lambda>\mu$ (as we did for PD), and we let $\rho=(\mu+\nu)/\lambda$. From [43] we have:
@@ 4.3 The relationship between $\varphi_{FD}$ and $\varphi_{PD}$
Moreover, numerical evaluations of $\varphi_{FD}(s)$ using eqn. (13) suggest that for $s\neq 0,1$ the value of $\varphi_{FD}(s)$ is strictly monotone decreasing with $\nu$ if the other parameters are fixed. This monotonicity has been established for the critical case where $\lambda=\mu+\nu$ (i.e., $\rho=1$), since in that case the expression for $\varphi_{FD}(s)$ (described above) becomes more tractable [47]. Moreover, this monotonicity (if true) would imply that the following inequality also holds (since $\varphi_{FD}=\varphi_{PD}$ when $\nu=0$).
$\varphi_{FD}(s)<\varphi_{PD}(s)$ (14)
provided that $\lambda>\mu$, $s\neq 0,1$ and $\nu>0$.
What can easily be established is that a variant of eqn. (14) holds, even on fixed trees (rather than just for birth-death trees). In this variation the temporal length of each edge (in the calculation of $PD$) is replaced the expected number of features that arise on the edge and which are present in at least one species at the present (prior to the FOB model). More precisely, suppose that $T$ is a phylogenetic $X$-tree $T$, with edge lengths $\ell=[\ell_{e}]$. To allow features to be present at the root of this tree, we add an additional stem edge above the root, and assume that no features are present at the start of this stem edge. For an edge $e$ of $T$ of length $\ell$, let $\ell^{\prime}_{e}$ be the expected number of features that (i) arise on edge $e$ and (ii) are present in at least one of the leaves of $T$, and let $\mathcal{S}$ be the set of species present after the FOB extinction event. The following inequality then holds (a proof is provided in the Appendix).
@@ 4.4 Concluding comments
If eqn. (14) is true generally, it would provide a relevant message for biodiversity theory. Namely, using PD as a proxy to measure FD will, in expectation, provide a systematic underestimate of the true proportion of feature diversity that is expected to be lost in an extinction event in the near future, at least for the simple models considered in this paper. It would therefore be interesting to investigate the degree to which such underestimation holds on real phylogenetic trees (perhaps within the framework of EvoHeritage [45]). A mathematical proof (or disproof) of eqn. (14) would also be helpful.
@@ 6 Appendix
where $\epsilon=(1-s)$ is the FOB extinction probability. Clearly, $\sum_{k=1}^{n_{e}}\phi^{T}_{e}(k)\epsilon^{k}\geq\sum_{k=1}^{n_{i}}\phi^{T}_{i}(k)\epsilon^{n_{e}}$, so:
$\frac{\mathbb{E}[FD_{(T,\ell)}(\mathcal{S})]}{\mathbb{E}[FD_{(T,\ell)}(X)]}\leq 1-\frac{\sum_{e}\tilde{\ell}_{e}\cdot(\sum_{k=1}^{n_{i}}(\phi^{T}_{e}(k)\epsilon^{n_{e}})}{\sum_{e}\tilde{\ell}_{e}\cdot(\sum_{k=1}^{n_{e}}\phi_{e}^{T}(k))}=1-\frac{\sum_{e}\ell^{\prime}_{e}\epsilon^{n_{e}}}{\sum_{e}\ell^{\prime}_{e}}=\frac{\mathbb{E}[PD_{(T,\ell^{\prime})}(\mathcal{S})]}{PD_{(T,\ell^{\prime})}(X)},$
where the first equality is because $\sum_{k=1}^{n_{i}}\phi^{T}_{i}(k)=1-\phi_{e}^{T}(0)$. Moreover, the inequality is strict when $s\neq 0,1$, $\nu>0$, and $T$ is not a star tree because in that case $\sum_{k=1}^{n_{e}}\phi^{T}_{i}(k)\epsilon^{k}>\sum_{k=1}^{n_{e}}\phi^{T}_{i}(k)\epsilon^{n_{e}}$ for any edge $e$ that is not a pendant edge of $T$. This completes the proof. $\Box$

### v3

@@ 4 Feature diversity (FD) - a generalisation of PD
It turns out that FD can be represented by PD on some tree when the collection $\mathcal{C}_{\mathbb{F}}$ of subsets $X_{f}=\{x\in X:f\in F_{x}\}$ is nested (i.e., the condition that $A,B\in\mathcal{C}_{\mathbb{F}}\Rightarrow A\cap B\in\{A,B,\emptyset\}$ holds), by results in [42] (Proposition 1 and Theorem 1). Moreover, since a nested collection of subsets of $X$ corresponds to a subset of the clusters of a phylogenetic tree, it is possible to describe the evolution of each character in $\mathbb{F}_{X}$ on that tree by a single-gain (of the feature on one edge $e$) and no subsequent loss (on any descendant edge of $e$). A slightly stronger statement (from [42]) is the following:
@@ 4.1 A stochastic model of FD on fixed trees
Figure 4: A simple example to illustrate the gain and loss of features on a tree and the FOB process (from [43]). Left: New features arise (indicated by $+$) and disappear (indicated by --) along edges of the tree. Each new features arises only once in the tree, and for convenience, no features are present at time 00. Five features are present among the leaves of the tree at time $t$: $\{\alpha,\beta,\gamma,\delta,\epsilon\}$. A rapid extinction event at the present then results in the loss of three leaves of the tree (denoted by $\dagger$). Right: Three features ($\alpha,\beta,\gamma$) remain among the leaves of the resulting tree. Thus, the ratio of surviving features to total features is $3/5=0.6$.
@@
  • a new feature arises (on any edge) at rate $r>0$ per unit time, and
@@
This simple model can be described by a classical (constant-birth, linear-death) stochastic process for which the following results hold (see e.g. [44], p. 461). Consider a directed path in the tree $T$ of length $L$ (the sum of the $\ell_{e}$ values on the path). Conditional on there being $k$ features present at the start of this path, let $N_{L}^{(k)}$ denote the number of features present at the end of this path. For $k=0$, $N_{L}^{(0)}$ has a Poisson distribution with mean $\frac{r}{\nu}(1-e^{-\nu L})$, and this mean is asymptotic to $rL$ as $L\rightarrow 0$ and to $r/\nu$ as $L\rightarrow\infty.$ More generally, if there are $k$ features present at the start of the path, then $N_{L}^{(k)}$ is the sum of two independent random variables: one distributed as $N_{L}^{(0)}$, the other has a Gamma distribution with parameters $(k,\nu L)$. In particular, as $L\rightarrow\infty$, $N_{L}^{(k)}$ converges to a Poisson distribution with mean $r/\nu$.
@@ 4.2 FD in birth–death trees with extinction at the present
In this setting, we are interested in determining the proportion of features at the tips of the tree that survive the FOB model.
@@
Let $\mathcal{T}_{t}^{s}$ denote the tree obtained from $\mathcal{T}_{t}$ by deleting all leaves at time $t$ that do not survive the FOB model. The number of species at time $t$ in $\mathcal{T}_{t}^{s}$ that have a copy of a particular feature that arose at some fixed time $t_{0}\in(0,t)$ in $\mathcal{T}_{t}$ is described exactly by the birth–death process $X_{t-t_{0}}$ with birth rate $\lambda$, death rate $\theta=\mu+\nu$ and extinction at the present (with survival probability $s$ for each individual). It follows that as $t$ becomes large, it becomes increasingly certain that none of the species at time $t$ in $\mathcal{T}_{t}$ will contain a given feature $f$ present at $t_{0}$ if $\lambda\leq\mu+\nu$, whereas if $\lambda>\mu+\nu$, there is a positive limiting probability that $f$ will be present in the extant leaves of the $\mathcal{T}_{t}$ (even after pruning under the FOB model).
@@
It is easily verified that $\varphi_{FD}(s)$ is an increasing concave function that satisfies $1\geq\varphi_{FD}(s)\geq s$ for all $s$. This is similar to the earlier story for PD (indeed when $\nu=0$, $\varphi_{FD}=\varphi_{PD}$)). However, in contrast to $\varphi_{PD}(s)$, there does not seem to be a simple explicit expression for $\varphi_{FD}(s)$ when $\nu>0$. Nevertheless, $\varphi_{FD}(s)$ can be expressed as a ratio of two integrals as follows. We will assume that $\lambda>\mu$ (as we did for PD), and we let $\rho=(\mu+\nu)/\lambda$. From [43] we have:
@@ 4.3 The relationship between $\varphi_{FD}$ and $\varphi_{PD}$
Numerical evaluations of $\varphi_{FD}(s)$ using eqn. (13) suggest that for $s\neq 0,1$ the value of $\varphi_{FD}(s)$ is strictly monotone decreasing with $\nu$ if the other parameters are fixed. This monotonicity has been formally established for the critical case where $\lambda=\mu+\nu$ (i.e., $\rho=1$), since in that case the expression for $\varphi_{FD}(s)$ (described above) becomes more tractable [47]. If monotonicity holds generally, this would imply that the following inequality also holds (since $\varphi_{FD}=\varphi_{PD}$ when $\nu=0$):
$\varphi_{FD}(s)<\varphi_{PD}(s),\mbox{ for all $\lambda>\mu$, $s\neq 0,1$ and $\nu>0$. }$ (14)
What can easily be established is that a variant of eqn. (14) holds, even on fixed trees (rather than just for birth-death trees). In this variation the temporal length of each edge (in the calculation of $PD$) is replaced by the expected number of features that arise on the edge and which are also present in at least one species at the present (prior to the FOB model). More precisely, suppose that $T$ is a phylogenetic $X$-tree $T$, with edge lengths $\ell=[\ell_{e}]$. To allow features to be present at the root of this tree, we add an additional stem edge above the root, and assume that no features are present at the start of this stem edge. For an edge $e$ of $T$ of length $\ell$, let $\ell^{\prime}_{e}$ be the expected number of features that arise on edge $e$ and, in addition, are present in at least one of the leaves of $T$, and let $\mathcal{S}$ be the set of species present after the FOB extinction event. The following inequality then holds (a proof is provided in the Appendix).
@@ 4.4 Concluding comments
If inequality (14) is true generally, it would provide a relevant message for biodiversity theory. Namely, using PD as a proxy to measure FD will, in expectation, provide a systematic underestimate of the true proportion of feature diversity that is expected to be lost in an extinction event in the near future, at least for the simple models considered in this paper. It would therefore be interesting to investigate the degree to which such underestimation holds on real phylogenetic trees (perhaps within the framework of EvoHeritage [45]). A mathematical proof (or disproof) of eqn. (14) would also be helpful.
@@ 6 Appendix
where $\epsilon=(1-s)$ is the FOB extinction probability. Now, since $k\leq n_{e}$ we have: $\mathbb{E}[FD_{(T,\ell)}(\mathcal{S})]\leq\sum_{e}\tilde{\ell_{e}}\cdot\left(\sum_{k=1}^{n_{e}}\phi^{T}_{e}(k)\right)(1-\epsilon^{n_{e}}),$ and so
$\frac{\mathbb{E}[FD_{(T,\ell)}(\mathcal{S})]}{\mathbb{E}[FD_{(T,\ell)}(X)]}\leq\frac{\sum_{e}\tilde{\ell}_{e}\cdot\left(\sum_{k=1}^{n_{e}}\phi^{T}_{e}(k)\right)(1-\epsilon^{n_{e}})}{\sum_{e}\tilde{\ell}_{e}\cdot\left(\sum_{k=1}^{n_{e}}\phi^{T}_{e}(k)\right)}.$ (15)
In addition, since $\sum_{k=1}^{n_{e}}\phi^{T}_{e}(k)=1-\phi_{e}^{T}(0)$, and $\ell_{e}^{\prime}=\tilde{\ell_{e}}\cdot(1-\phi_{e}^{T}(0))$ (by definition), Inequality (15) can be rewritten as
$\frac{\mathbb{E}[FD_{(T,\ell)}(\mathcal{S})]}{\mathbb{E}[FD_{(T,\ell)}(X)]}\leq\frac{\sum_{e}\ell^{\prime}_{e}(1-\epsilon^{n_{e}})}{\sum_{e}\ell^{\prime}_{e}}=\frac{\mathbb{E}[PD_{(T,\ell^{\prime})}(\mathcal{S})]}{PD_{(T,\ell^{\prime})}(X)},$
which establishes the claimed inequality. Moreover, the inequality is strict when $s\neq 0,1$, $\nu>0$, and $T$ is not a star tree because in that case $\sum_{k=1}^{n_{e}}\phi^{T}_{e}(k)\epsilon^{k}>\sum_{k=1}^{n_{e}}\phi^{T}_{e}(k)\epsilon^{n_{e}}$ for any edge $e$ that is not a pendant edge of $T$. $\Box$